OpenStudy (anonymous):

I don't understand how f(x)=(tan(x)-2)/sec(x) gives f'(x)= 2sin(x)cos(x). Can someone show me how this was done. Please and thank you.

zepdrix (zepdrix):

Do you know your tangent and secant derivatives?

OpenStudy (anonymous):

tangent is sec^2x and secant is tanxsecx

OpenStudy (anonymous):

(f'g-g'f)/g^2 so (sec^2x(secx)-tanxsecx(tanx))/sec^2x right?

zepdrix (zepdrix):

Oh oh oh I see the problem.

zepdrix (zepdrix):

Yah there was a little boo boo in your derivative.

zepdrix (zepdrix):

\[\Large\rm =\frac{\color{royalblue}{\left(\tan x-2\right)'}\sec x-(\tan x-2)\color{royalblue}{\left(\sec x\right)'}}{\sec^2x}\] \[\Large\rm =\frac{\color{orangered}{\left(sec^2x\right)}\sec x-(\tan x-2)\color{orangered}{\left(\sec x tan x\right)}}{\sec^2x}\]Do you see how this is a little different than what you wrote?

zepdrix (zepdrix):

You forgot the -2 from the tanx-2 in the second term.

OpenStudy (anonymous):

Your right. Oups

zepdrix (zepdrix):

Still factoring out that secant is a good idea, and cancelling out if able,\[\Large\rm =\frac{\cancel{\sec x}\left(\sec^2x-(\tan x-2)\tan x\right)}{\sec^{\cancel{2}1}x}\]

zepdrix (zepdrix):

Which leaves us with this: \[\Large\rm =\frac{\color{royalblue}{\sec^2x-\tan^2x}+2\tan x}{\sec x}\]

zepdrix (zepdrix):

And we used our Pythagorean Identity to show that this blue part is simply 1,\[\Large\rm =\frac{\color{royalblue}{1}+2\tan x}{\sec x}\]

zepdrix (zepdrix):

Mmmm what do you think? :d Able to simplify it the rest of the way from there? Confused by any of those steps?

OpenStudy (anonymous):

no I am following. Makes more sense now that the 2 is there.

OpenStudy (anonymous):

Thank you very much for your help

zepdrix (zepdrix):

No problem!